OpenStudy (anonymous):

A ladder leans against a house at a 60° angle to the ground. If the ladder extends to a length of 158 inches, what is the height of the house rounded to the nearest hundredth of an inch? Type the numeric answer only in the box below.

OpenStudy (anonymous):

ratio of 30 - 60 - 90 triangle are \(1:\sqrt{3}:2\)

OpenStudy (anonymous):

take \[158\times \frac{\sqrt{3}}{2}\]

OpenStudy (anonymous):

136.83?